Find both experimental error and the percent error.
QUESTION:
experimental value: 2.4 ; known value: 2.5

Answer by Theo(13342)

here's a reference that discusses absolute error and relative error.
http://www.regentsprep.org/regents/math/algebra/am3/LError.htm

not sure if you're dealing with absolute error.
that is the absolute value of the error.

it doesn't really tell you if you were below or above the desired measurement.

if that's a consideration, then you would need to indicate the direction of the error as well.

your experimental value is 2.4 and your known value is 2.5.

the absolute value of the difference is |2.4 - 2.5| = |-.1| = .1

the relative difference is the absolute value of the difference divided by the known value.

this would be .1 / 2.5 = .04 * 100 = 4%.

so you're talking an absolute difference of .1 and a relative difference of 4%.

if the direction of the difference is important, then you would need to say that the absolute difference is .1 below the known value and the relative difference is 4% of the known value.